OpenStudy (anonymous):

Two polygons are congruent and the perimeter of the first polygon is 38 cm. If the sides of the second polygon are consecutive integers (x, x+1, x+2, x+3, etc.), what value of x makes the polygons into congruent quadrilaterals? Use x as the smallest side.

OpenStudy (anonymous):

38 = x + x + 1 + x + 2 + x + 3 38 = 4x + 6 32 = 4x x = 8 If the two polygons are congruent, that means all parts of it are congruent as well. This includes area, perimeter, and side lengths. Plus, it asks congruent QUADRILATERALS so you only use 4 of the consecutive sides.

OpenStudy (anonymous):

would x= 8 by any chance?

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

lol posted that just as ur response came in thank you

OpenStudy (anonymous):

Yes you are right. lol You're welcome :D

OpenStudy (anonymous):

mind if i ask you one more

OpenStudy (anonymous):

Sure Go ahead!

OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 225:36. What is the ratio of their corresponding sides?

OpenStudy (anonymous):

would it be 5:2?

OpenStudy (anonymous):

Unfortunately no. Hold on and I'll explain the concept.

OpenStudy (anonymous):

First you have to understand the concept that similar polygons aren't just similar in terms of side lengths. Their areas also have a special ratio. The ratio of side lengths is represented as a:b. The ratio of areas is a^2:b^2. So, if a^2:b^2 = 225:36, then a:b = 15:6 or 5:2 because to get a:b, you find the square root. The square root of 225 is 15 and the square root of 36 is 6.

OpenStudy (anonymous):

lol so i was right?

OpenStudy (anonymous):

O_O lol. I didn't see that you simplified it. -_- That's what I get for doing it in my head...

OpenStudy (anonymous):

do you mind helping me check a few more lol dont want to get a bad score on this

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

If a picture with a width of 4 inches is made larger using a ratio of similarity of 2:5, what is the new width of the picture?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

seems obvious but 2 simple to be right lol

OpenStudy (anonymous):

No. It's 10. You set up a proportion. 2/5 = 4/x THis turns to 20 = 2x. x = 10

OpenStudy (anonymous):

If you think logically, you just doubled 2 to get 4. So, you would just multiply 5 by 2 to get 10 if you want to do it sensibly.

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

last 1 xD

OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF and has a ratio of similarity of 6:1. If side ED measures 12 mm, how long is side BA?

OpenStudy (anonymous):

BA is 2.

OpenStudy (anonymous):

Wait a sec. Nevermind I goofed

OpenStudy (anonymous):

BA is 72.

OpenStudy (anonymous):

thank you xD submitted the exam and got a perfect score

OpenStudy (anonymous):

You're Welcome

OpenStudy (anonymous):

virtual school would be a pain if not for this site

OpenStudy (anonymous):

Congratulations! :)

OpenStudy (anonymous):

Yeah. I know what it feels like. I'm finishing Algebra II online around now.

OpenStudy (anonymous):

but you're good at math so to you its a matter of when you feel like doing.

OpenStudy (anonymous):

doing it*

OpenStudy (anonymous):

Kind of, but sometimes it's a pain showing your work on a computer. Especially with all of the special symbols.

OpenStudy (anonymous):

well if you're bored let me know i have more questions to ask moving onto the next quiz pretty much on the same material

OpenStudy (anonymous):

youre help would make this go by super fast

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

Sure. I'll help.

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

k the next question is triangle bcd is equal or similar to triangle fgh if side cd is equal to side gh and side bd is equal or similar to side fh then which angle is used to show the sas postulate

OpenStudy (anonymous):

Angle D. Because the angle is included.

OpenStudy (anonymous):

these are the options a)angle bcd and angle fgh b)angle bdc and angle fhg c)angle cbd and angle gfh d)angle dcb and angle hgf

OpenStudy (anonymous):

Choice B. I believe

OpenStudy (anonymous):

k ty that was the only annoying 1

OpenStudy (anonymous):

heres the next one

OpenStudy (anonymous):

What is the area of a regular pentagon with a side of 7.2 and an apothem of 6?

OpenStudy (anonymous):

Ok. THis one is easier than you'd think. A = 1/2aP where a is the apothem and P is the perimeter. A = 108

OpenStudy (anonymous):

how do you remember all these formulas lol

OpenStudy (anonymous):

lol. You just learn to remember them. I'm in math team at school, so I kinda have to remember in order to do well because I also have no calculator. That's why I can do this stuff in my head too. :P

OpenStudy (anonymous):